OpenStudy (anonymous):

Can someone please help me? I really don't get this: 2. Given the parent function f(x) = x2, describe the translation of the function: f (x) = 1/4 (x-6)^2 + 8 A. 6 units to the left, 8 units up B. 6 units to the left, 8 units down C. 6 units to the right, 8 units up D. 6 units the right, 8 units down

OpenStudy (anonymous):

Given the parent function f(x) = x2, describe the translation of the function f(x)=1/4(x-6)^2+8? a. 6 units to the left, 8 units up b. 6 units to the left, 8 units down c. 6 units to the right, 8 units up [Answer] d. 6 units the right, 8 units down Good luck!

OpenStudy (anonymous):

The answer has been given above. By way of explanation, the +8 means that all the values of f(x) (the 'y' values) will be increased by 8 and so the graph moves 'up'. Now, consider y = x^2 and y = 1/4 (x-6)^2 when y is 0 then x is 0 and 6, respectively so the graph is moved to the 'right' (from 0 to 6)

OpenStudy (anonymous):

Some people see things algebraically, others geometrically, mess about with different ways to find one that feels natural to you.